### v1

# Feature Engineering is Not Dead: Reviving Classical Machine Learning with Entropy, HOG, and LBP Feature Fusion for Image Classification

Abstract
Feature engineering continues to play a critical role in image classification, particularly when interpretability and computational efficiency are prioritized over deep learning models with millions of parameters. In this study, we revisit classical machine learning based image classification through a novel approach centered on Permutation Entropy (PE), a robust and computationally lightweight measure traditionally used in time series analysis but rarely applied to image data. We extend PE to two-dimensional images and propose a multiscale, multi-orientation entropy-based feature extraction approach that characterizes spatial order and complexity along rows, columns, diagonals, anti-diagonals, and local patches of the image. To enhance the discriminatory power of the entropy features, we integrate two classic image descriptors: the Histogram of Oriented Gradients (HOG) to capture shape and edge structure, and Local Binary Patterns (LBP) to encode micro-texture of an image. The resulting hand-crafted feature set, comprising of 780 dimensions, is used to train Support Vector Machine (SVM) classifiers optimized through grid search. The proposed approach is evaluated on multiple benchmark datasets, including Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10, where it delivers competitive classification performance without relying on deep architectures. Our results demonstrate that the fusion of PE with HOG and LBP provides a compact, interpretable, and effective alternative to computationally expensive and limited interpretable deep learning models. This shows a potential of entropy-based descriptors in image classification and contributes a lightweight and generalizable solution to interpretable machine learning in image classification and computer vision.
Index Terms:
Permutation Entropy, Image Classification, Feature Fusion, Classical Machine Learning, Interpretable Models, Histogram of Oriented Gradients (HOG), Local Binary Patterns (LBP), Support Vector Machines (SVM), Benchmark Datasets (e.g., Fashion-MNIST, CIFAR-10).
I Introduction
Permutation entropy (PE), introduced by Bandt and Pompe [1], is a symbolic time-series complexity measure that captures the diversity of ordinal patterns in a signal. Unlike conventional entropy metrics such as Shannon or Rényi entropy [2, 3], which are derived from amplitude-based probability distributions, PE encodes local temporal dynamics by analyzing the relative ordering of values within sliding windows. Due to its robustness to noise, computational simplicity and model-free formulation, PE has found widespread application in physiological signal analysis, including EEG-based seizure detection [4], heart rate variability [5], and sleep stage classification [6].
Despite its success in 1D temporal domains, PE has rarely been explored in image classification. A few works have used PE in medical imaging scenarios—such as patch-wise entropy for CT [7] and fMRI [8] analysis—but these have been limited in scope, focusing on narrow domains or local feature statistics. To the best of our knowledge, PE has never been employed as a primary, global feature descriptor for natural image classification.
This gap is particularly notable given the computational cost and complexity associated with modern deep learning approaches. Convolutional neural networks (CNNs) such as VGG [9], ResNet [11], and MobileNet [12] dominate datasets like Fashion-MNIST [13], achieving 93–94% classification accuracy. However, these models come at the expense of millions of trainable parameters, extensive training resources, and limited interpretability. For example, ResNet-18 contains more than 11 million parameters [11], and even lightweight models such as MobileNet require substantial GPU memory and training time. Furthermore, the black-box nature of deep learning impedes transparency and scientific understanding.
In this work, we present a novel, interpretable, and computationally lightweight approach that uses PE for full-image classification. The PE features are combined with two complementary image descriptors, the Histogram of Oriented Gradients (HOG) [14] and the Local Binary Patterns (LBP) [15]. HOG captures shape and edge structure of an image, useful for detecting object outlines and silhouettes, whereas LBP captures micro-texture such as spots, flat areas, edges and corners, and good at capturing fine repetitive patterns. We combine PE, HOG and LBP features to form a compact, handcrafted feature vector of n-dimensions. We use these handcrafted feature vectors and employ Support Vector Machines (SVM) [16], a classical machine learning method, for image classification work across multiple benchmark datasets such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10. In particular, we achieve up to 91.15% test accuracy on Fashion-MNIST benchmark dataset. All previous test accuracy results [17] in the Fashion-MNIST benchmark dataset, that used scikit-learn library and evaluated 129 classifiers with various parameter configurations (excluding deep learning methods), reported a maximum accuracy of 89.7% using SVM with parameters C=10, kernel=poly [17]. In contrast, our method achieves an accuracy of 91.15% for the first time using purely classical machine learning techniques applied on hand-crafted features extracted from the Fashion-MNIST dataset. We also apply this method on other benchmark datasets (KMNIST, EMNIST, and CIFAR-10) and show competitive performance.
The rest of the paper is organized as follows: In Section II, we present a detailed review of permutation entropy; Section III discusses the classical image descriptors HOG and LBP employed in this work; Section IV describes our feature-extraction pipeline in detail; Section V briefly introduces classical SVM-based machine learning; Section VI explains our rationale for selecting the benchmark datasets and provides their descriptions; Section VII reports the results obtained using our novel feature-extraction pipeline and compares them with previous studies; finally, Section VIII concludes the paper and outlines future implications.
II Permutation Entropy (PE)
Shannon entropy is a foundational concept in information theory that quantifies the average amount of uncertainty or information contained in a probability distribution. For a discrete random variable $X$ with possible outcomes $\{x_{1},x_{2},\dots,x_{n}\}$ and corresponding probabilities $\{p_{1},p_{2},\dots,p_{n}\}$, the Shannon entropy is defined as:
$H(X)=-\sum_{i=1}^{n}p_{i}\log p_{i}.$ (1)
Entropy reaches its maximum when all outcomes are equally likely (i.e., the distribution is uniform), and is lower when the distribution is more predictable or concentrated.
Permutation Entropy is a specialized form of entropy introduced by Bandt and Pompe [1] to measure the complexity of time series data. Unlike Shannon entropy, which requires an explicit probability distribution over symbols or values, PE evaluates the entropy of the ordinal patterns — that is, the relative ordering of values within segments of the time series.
Figure 1: Visualization of PE from a time-series data: (a) a sample time series data; and (b) corresponding normalized PE with embedding dimension d = 3 and time delay $\tau$ =1. Note that in high irregularity domain PE is near 1 and in high regularity or predictable domain PE is near 0.
To understand PE, let’s consider a scalar time series $\{x_{t}\}_{t=1}^{N}$. To extract ordinal patterns, we first embed the series in a higher-dimensional space using parameters embedding dimension $d\geq 2$ (length of each pattern) and time delay $\tau\geq 1$ (step between points in a pattern). This results in a sequence of delay vectors:
$\mathbf{X}_{t}=($ $x_{t},x_{t+\tau},x_{t+2\tau},\dots,$
$x_{t+(d-1)\tau}),\quad t=1,2,\dots,N-(d-1)\tau.$ (2)
For each vector $\mathbf{X}_{t}$, we determine its ordinal pattern — the permutation $\pi\in S_{d}$ that sorts the elements in ascending order:
$x_{t+\pi_{1}\tau}\leq x_{t+\pi_{2}\tau}\leq\cdots\leq x_{t+\pi_{d}\tau}.$
If ties occur, they are resolved by the order of appearance. Next, we compute the relative frequency $P(\pi)$ of each permutation $\pi$ across the time series. The PE is then defined as the Shannon entropy of this distribution:
$H_{d}=-\sum_{\pi\in S_{d}}P(\pi)\log P(\pi),$ (3)
where $S_{d}$ is the set of all $d!$ possible permutations of order $d$. To enable comparison across different embedding dimensions, it is often normalized:
$h_{d}=\frac{H_{d}}{\log(d!)}\in[0,1].$ (4)
For example, let us consider a short time series: $\{x_{t}\}=\{4,7,9,10,6\}$, with $d=3$ and $\tau=1$. The delay vectors are:
$\mathbf{X}_{1}$ $=(4,7,9)\rightarrow\pi=(1,2,3),$
$\mathbf{X}_{2}$ $=(7,9,10)\rightarrow\pi=(1,2,3),$
$\mathbf{X}_{3}$ $=(9,10,6)\rightarrow\pi=(3,1,2).$
From these, the ordinal patterns $(1,2,3)$ and $(3,1,2)$ occur with frequencies $2/3$ and $1/3$, respectively. The entropy is then:
$H_{3}=-\left(\frac{2}{3}\log\frac{2}{3}+\frac{1}{3}\log\frac{1}{3}\right)\approx 0.918\text{ bits},$
$h_{3}=\frac{H_{3}}{\log(6)}\approx\frac{0.918}{\log(6)}\approx 0.51.$
A low PE (near 0) indicates high regularity or predictability—common in periodic or deterministic signals. In contrasts a high PE (near 1) suggests randomness or high complexity—typical of stochastic or chaotic processes (Fig. 1).
In the context of image analysis, 2D grayscale image patches can be flattened into 1D sequences and analyzed using PE to quantify local structural complexity. This is particularly useful in medical imaging, texture classification, or anomaly detection, where small variations in spatial ordering can signal important features. Unlike pixel intensity-based methods, PE is sensitive to the shape and variation patterns, making it suitable for evaluating datasets with rich intra-class texture or fine-grained structural differences.
III Traditional Feature Descriptors: HOG and LBP
III-A Histogram of Oriented Gradients (HOG)
HOG is a widely used feature descriptor for object detection and shape-based analysis, originally proposed by Dalal and Triggs for pedestrian detection [14]. The core idea is to characterize local object appearance and shape by the distribution of edge orientations (gradients) (Fig. 2). Given a grayscale image $I(x,y)$, the first step is to compute the gradient at each pixel using finite difference filters:
$G_{x}=I(x+1,y)-I(x-1,y),\quad G_{y}=I(x,y+1)-I(x,y-1).$ (5)
From these gradients, the magnitude $M(x,y)$ and orientation $\theta(x,y)$ are computed as:
$M(x,y)=\sqrt{G_{x}^{2}+G_{y}^{2}},\quad\theta(x,y)=\arctan 2(G_{y},G_{x}).$ (6)
Figure 2: Visualization of HOG from an image: (a) reference image (source: scikit-image); (b) corresponding HOG; (c) Schematic visualization of 9 HOG Orientation Bins (0°–180°, unsigned; arrows = gradient directions). Note that a gradient at 90° (arrow pointing straight up) comes from a horizontal edge, while a gradient at 0°/180° (arrow along the x-axis) comes from a vertical edge; (d) plot of global HOG orientation histogram (9 orientation bins, 0°–180°) in which each bar aggregates the total gradient magnitude for orientations falling into its 20° bin.
The image is divided into small connected regions called cells (e.g., $8\times 8$ pixels), and a histogram of gradient directions is computed within each cell, typically with 9 orientation bins spanning $0^{\circ}$ to $180^{\circ}$ (unsigned) (Fig. 2). Each pixel contributes to the histogram proportionally to its gradient magnitude.
To enhance robustness to illumination and contrast variations, the histograms are normalized across larger regions called blocks (e.g., $2\times 2$ cells), using $L_{2}$ or $L_{1}$-norm normalization:
$v\leftarrow\frac{v}{\sqrt{\|v\|_{2}^{2}+\epsilon^{2}}},$ (7)
where $v$ is the concatenated vector of histograms in a block and $\epsilon$ is a small constant to avoid division by zero.
HOG features are translation-invariant at the cell level and capture edge and contour information that is highly discriminative for structured objects.
III-B Local Binary Pattern (LBP)
LBP is a texture descriptor that encodes local spatial patterns by thresholding the neighborhood of each pixel. It was first introduced by Ojala et al. [15] as a powerful and simple operator for texture classification.
For a pixel at position $(x,y)$, and a circular neighborhood of $P$ pixels at radius $R$, the LBP code is computed as:
$\text{LBP}_{P,R}(x,y)=\sum_{p=0}^{P-1}s(g_{p}-g_{c})\cdot 2^{p},$ (8)
Here, $g_{c}$ denotes the gray value of the center pixel and $g_{p}$ denotes the gray value of the $p$-th neighbor (interpolated if it does not lie exactly on the pixel grid). The function
$s(z)=\begin{cases}1,&\text{if }z\geq 0,\\ 0,&\text{otherwise,}\end{cases}$
takes the value 1 when $z\geq 0$ and 0 otherwise.
This results in a binary number (e.g., 8-bit for $P=8$) representing the local texture pattern. A histogram of LBP codes is then computed over local image regions and concatenated to form the final descriptor.
LBP is highly discriminative for micro-patterns such as edges, corners, and spots, and it is invariant to monotonic gray-scale transformations, making it robust to illumination changes (Fig. 3).
Figure 3: Visualization of LBP from an image: (a) reference image (source: scikit-image); (b) corresponding LBP.
Combined Use in Feature Fusion: When used together, HOG and LBP provide complementary information: HOG captures global structure and shape, while LBP encodes fine-scale local texture. Note that although HOG is more of an edge/shape descriptor, but its dense‐grid histogram approach can incidentally pick up some texture cues (Fig. 2). In contrasts, LBP is explicitly a texture descriptor (Fig. 3). Feature fusion of HOG and LBP has been shown to improve classification performance in various tasks, including face recognition, scene classification, and biomedical imaging [19].
IV The Feature Extraction Pipeline: The Fusion of Complexity, Correlation, Shape and Texture
IV-A Feature Extraction Process and Dimensions
In this work, each image from standard benchmark datasets (such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10) is transformed into a handcrafted feature vector composed of entropy-based, correlation-based, shape and texture-based image descriptors. Although dimensionality may vary slightly between datasets due to image size or content, the core extraction mechanisms remain consistent (Fig. 4). In the following, we provide a general overview of the feature types and the methods used for their extraction.
Row-wise PE: Each row in the image is treated as a 1D signal. For each row vector $\mathbf{r}$, both the forward $PE(\mathbf{r})$ and reversed permutation entropy $PE(\mathbf{r}^{\text{rev}})$ are calculated, which means that each row of the image is treated as a one-dimensional signal: for a given row vector $\mathbf{r}$, the sequence of pixel values is interpreted first in the forward (left-to-right) direction and then in the reversed (right-to-left) direction, and the permutation entropy is computed for both resulting time series. Finally, the final geometric mean PE is calculated as follows:
$PE_{\text{row}}^{GM}=\sqrt{PE(\mathbf{r})\cdot PE(\mathbf{r}^{\text{rev}})}$
yielding one feature per row. We do so for all possible rows.
Column-wise PE: Analogously to the row-wise case, for each column vector $\mathbf{c}$ (note that this vector $\mathbf{c}$ can be treated as 1D signal/time series data) we compute the forward and reversed permutation entropies and combine them using geometric mean as follows:
$PE_{\text{col}}^{GM}=\sqrt{PE(\mathbf{c})\cdot PE(\mathbf{c}^{\text{rev}})}$
yielding one feature per column.
Diagonal PE: We simplify the diagonal feature extraction as follows. Starting with the main diagonal (top-left to bottom-right), we also take the $K$ diagonals above it and the $K$ diagonals below it—so in total $12K+$ 1 diagonals. For each diagonal line, we read its pixel values in order to form a one-dimensional sequence, then compute permutation entropy on that sequence and again on its reversed sequence. By taking the geometric mean of these two entropy values, we collapse each diagonal into a single feature . This reflects how complex the pixel arrangement is along that slanted direction.
Thus, for each diagonal $d_{k}$ we have:
$PE_{\text{diag}}(k)=\frac{PE(d_{k})+PE(d_{k}^{\text{rev}})}{2}$
Anti-Diagonal PE: We handle anti-diagonals in exactly the same way. An anti-diagonal is simply a line of pixels running from the top-right corner down to the bottom-left corner (and any lines parallel to it). Concretely, we take the main anti-diagonal (top-right to bottom-left) and its $K$ offsets on either side, thus again yielding $12K+$ 1 lines—read each line’s pixel values into a one-dimensional sequence, compute permutation entropy on that sequence and on its reverse, then combine them via the geometric mean. This produces one scalar feature per anti-diagonal that, like the diagonal features, captures the complexity of pixel arrangements along that slanted orientation.
Thus, for each anti-diagonal $ad_{k}$ we have:
$PE_{\text{anti-diag}}(k)=\frac{PE(ad_{k})+PE(ad_{k}^{\text{rev}})}{2}$
Patch-wise PE: The image is divided into overlapping patches (e.g., $4\times 4$ with stride 2), and the PE is calculated for each patch:
$PE_{\text{patch}}(i)=PE(p_{i})$
Row and Column Correlation: Pearson correlation coefficients are computed between each pair of adjacent rows and adjacent columns:
$\rho_{i,i+1}^{\text{row}}=\text{corr}(\mathbf{r}_{i},\mathbf{r}_{i+1}),\quad\rho_{j,j+1}^{\text{col}}=\text{corr}(\mathbf{c}_{j},\mathbf{c}_{j+1})$
HOG features: HOG features are computed using standard cell/block configurations (e.g., $4\times 4$ pixels per cell) and a fixed number of orientation bins.
LBP features: LBP descriptors are extracted with a fixed radius and number of sampling points (e.g., $R=2$, $P=16$), resulting in a histogram of uniform patterns.
Figure 4: (a) Schematic overview of the proposed workflow pipeline. For each input image we extract permutation-entropy (PE) based features—computed along rows, columns, the diagonal and anti-diagonals, and on overlapping local patches. We combine them with row-/column-correlation based features, histogram of oriented gradients (HOG) and local binary pattern (LBP) image descriptors. The four feature sets are concatenated into a single feature vector and classified with a support vector machine (SVM). (b) Detailed illustration of the PE-based feature-extraction steps on a (6 × 6) toy image. PE is first evaluated for every row and column, followed by the diagonal (top-left → bottom-right) and anti-diagonal (top-right → bottom-left) directions. Note that some off-centre diagonals and anti-diagonals (towards corner) PEs are not computed because their reduced length may yield unreliable PE features or statistics. Finally, we compute patch-wise PE with (4 × 4) patches and with a stride length of 2 pixels. See Section IV for a complete description of the feature-extraction pipeline.
TABLE I: Toy $6\times 6$ image used to illustrate feature extraction. All pixel values $a_{ij}$ of the image are shown in the table
1 2 3 4 5 6
1 $a_{11}$ $a_{12}$ $a_{13}$ $a_{14}$ $a_{15}$ $a_{16}$
2 $a_{21}$ $a_{22}$ $a_{23}$ $a_{24}$ $a_{25}$ $a_{26}$
3 $a_{31}$ $a_{32}$ $a_{33}$ $a_{34}$ $a_{35}$ $a_{36}$
4 $a_{41}$ $a_{42}$ $a_{43}$ $a_{44}$ $a_{45}$ $a_{46}$
5 $a_{51}$ $a_{52}$ $a_{53}$ $a_{54}$ $a_{55}$ $a_{56}$
6 $a_{61}$ $a_{62}$ $a_{63}$ $a_{64}$ $a_{65}$ $a_{66}$
To illustrate the feature extraction process, we use a $6\times 6$ toy image whose pixel values $a_{ij}$ are listed in Table I. We calculate the Row-wise PE, for each row $i$ ($i=1,\dots,6$), and form the sequence
$r_{i}=\bigl{(}a_{i1},\,a_{i2},\,\dots,\,a_{i6}\bigr{)},$
and compute its permutation entropy in the forward direction,
$PE\bigl{(}r_{i}\bigr{)},$
and in the reversed direction,
$PE\bigl{(}r_{i}^{\mathrm{rev}}\bigr{)},\quad r_{i}^{\mathrm{rev}}=(a_{i6},\,a_{i5},\,\dots,\,a_{i1}).$
Finally, the geometric‐mean PE feature for row $i$ is
$PE_{\mathrm{row},\,i}^{GM}=\sqrt{\,PE\bigl{(}r_{i}\bigr{)}\times PE\bigl{(}r_{i}^{\mathrm{rev}}\bigr{)}\,}.$
Applying this to $i=1,\ldots,6$ produces the six row‐wise features $\{PE_{\mathrm{row},1}^{GM},\,PE_{\mathrm{row},2}^{GM},\,\dots,\,PE_{\mathrm{row},6}^{GM}\}$.
In case of Column-wise PE, for each column $j=1,\dots,6$, define
$c_{j}=\bigl{(}a_{1j},\,a_{2j},\,\dots,\,a_{6j}\bigr{)},c_{j}^{\mathrm{rev}}=(a_{6j},\,a_{5j},\,\dots,\,a_{1j}).$
Compute $PE(c_{j})$ and $PE(c_{j}^{\mathrm{rev}})$, then form
$PE_{\mathrm{col},\,j}^{GM}=\sqrt{\,PE(c_{j})\,\times\,PE(c_{j}^{\mathrm{rev}})\,}\,,\quad j=1,\dots,6,$
[…]
Let us now move to Diagonal-wise PE with $K=2$. For each offset $k=-2,-1,0,1,2$ define the diagonal
[…]
$d_{k}^{\mathrm{rev}}$ $=\bigl{(}a_{q,\;q+k}\;\bigm{|}\;q\text{ in reverse order}\bigr{)}.$
$d_{-2}$ $=(a_{31},\,a_{42}),$ $d_{-2}^{\mathrm{rev}}$ $=(a_{42},\,a_{31}),$
$d_{-1}$ $=(a_{21},\,a_{32},\,a_{43},\,a_{54}),$ $d_{-1}^{\mathrm{rev}}$ $=(a_{54},\,a_{43},\,a_{32},\,a_{21}),$
$d_{0}$ $=(a_{11},\,a_{22},\,...,\,a_{66}),$ $d_{0}^{\mathrm{rev}}$ $=(a_{66},\,a_{55},\,...,\,a_{11}),$
$d_{1}$ $=(a_{12},\,a_{23},\,a_{34},\,a_{45},\,a_{56}),$ $d_{1}^{\mathrm{rev}}$ $=(a_{56},\,a_{45},\,a_{34},\,a_{23},\,a_{12}),$
$d_{2}$ $=(a_{13},\,a_{24},\,a_{35},\,a_{46}),$ $d_{2}^{\mathrm{rev}}$ $=(a_{46},\,a_{35},\,a_{24},\,a_{13}).$
Compute $PE(d_{k})$ and $PE(d_{k}^{\mathrm{rev}})$ for five diagonal features.
$PE_{\mathrm{diag}}^{GM}(k)=\sqrt{\,PE(d_{k})\,\times\,PE(d_{k}^{\mathrm{rev}})\,}\,,\quad k=-2,\ldots,2,$
Similarly, for each $k=-2,-1,0,1,2$ define the anti‐diagonal
[…]
$ad_{k}^{\mathrm{rev}}=\bigl{(}a_{q,\;7-q+k}\;\bigm{|}\;q\text{ in reverse order}\bigr{)}.$
Compute $PE(ad_{k})$ and $PE(ad_{k}^{\mathrm{rev}})$, then for five anti‐diagonal features.
$PE_{\mathrm{anti}}^{GM}(k)=\sqrt{\,PE(ad_{k})\,\times\,PE(ad_{k}^{\mathrm{rev}})\,}\,,\quad k=-2,\ldots,2,\\ \$
V Support Vector Machine (SVM) for image classification
SVM [16] are a robust class of supervised learning models that seek an optimal separating hyperplane by solving a regularized hinge loss minimization problem:
$\min_{\mathbf{w},b}\frac{1}{2}\|\mathbf{w}\|^{2}+C\sum_{i=1}^{n}\max(0,1-y_{i}(\mathbf{w}^{T}\phi(\mathbf{x}_{i})+b)),$ (9)
where, $\mathbf{w}\in\mathbb{R}^{d}$ — The weight vector, normal to the separating hyperplane in feature space.
$b\in\mathbb{R}$ — The bias (intercept) term, which shifts the hyperplane away from the origin.
$\|\mathbf{w}\|^{2}$ — The squared Euclidean norm of $\mathbf{w}$, penalizing large weights to encourage a wide margin.
$C>0$ — The regularization parameter that balances margin maximization (small $\|\mathbf{w}\|$) against misclassification errors; larger $C$ increases the penalty on hinge‐loss violations.
$\phi:\mathbb{R}^{p}\to\mathcal{H}$ — An (often implicit) feature map into a (possibly high‐ or infinite‐dimensional) Hilbert space $\mathcal{H}$. Kernels allow evaluation of inner products in $\mathcal{H}$ without explicit computation of $\phi$.
$y_{i}\in\{+1,-1\}$ — The true class label of example $i$.
$\mathbf{x}_{i}\in\mathbb{R}^{p}$ — The input feature vector of example $i$.
[…]
In this work, we employ SVM not in their traditional raw pixel input setting [18], but in conjunction with a novel handcrafted feature extraction approach based on permutation entropy. To further enhance discriminative power, these entropy-based descriptors are combined with classical texture features: histogram of oriented gradients and local binary patterns. The result is a compact 780-dimensional feature vector that is interpretable, low-dimensional, and well-suited for classical machine learning models such as SVM.
Using this feature representation, we train SVM on four benchmark datasets: Fashion-MNIST, KMNIST, EMNIST (Balanced), and CIFAR-10. The use of handcrafted features allows SVM to operate effectively in a reduced-dimensional space, avoiding the scalability bottlenecks associated with high-dimensional kernel computations.
VI Benchmark Datasets: Fashion-MNIST, KMNIST, EMNIST, and CIFAR -10
VI-A Fashion-MNIST Dataset
Fashion-MNIST is a benchmark dataset consisting of 70,000 grayscale images, each with a resolution of $28\times 28$ pixels, categorized into 10 classes of clothing items: T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots [13]. The dataset is evenly split into 60,000 training images and 10,000 test images, with each class uniformly represented.
Despite its relatively small image size and lack of color information, Fashion-MNIST presents a meaningful challenge due to high intra-class variability and subtle inter-class differences. For example, distinguishing between shirts and T-shirts or between sneakers and ankle boots often requires sensitivity to texture, contour, and fine-grained shape differences. The images also exhibit varied orientations, partial occlusions, and diverse styles across classes.
These attributes make Fashion-MNIST a valuable resource for evaluating feature extraction methods that aim to capture structural and spatial patterns. Its grayscale format ensures that classification performance is driven primarily by shape and texture, rather than color, which aligns well with entropy-based approaches such as permutation entropy [1].
VI-B KMNIST Dataset
KMNIST (Kuzushiji-MNIST) is a drop-in replacement for MNIST, created to encourage research into Japanese character recognition [20]. It contains 70,000 $28\times 28$ grayscale images of cursive Japanese (Kuzushiji) characters from 10 classes. Unlike Latin characters in MNIST or fashion items in Fashion-MNIST, the cursive nature of Kuzushiji introduces unique challenges such as high intra-class variability, stroke overlap, and complex calligraphic patterns.
KMNIST’s higher visual complexity, despite its small size and grayscale format, presents an ideal middle ground between simplistic datasets like MNIST and natural image datasets like CIFAR-10. Its emphasis on subtle spatial structures makes it particularly compatible with entropy-based feature extraction approaches, which aim to quantify local complexity and variability within pixel arrangements.
VI-C EMNIST Dataset
EMNIST (Extended MNIST) expands upon the original MNIST dataset by including handwritten letters in addition to digits [22]. It consists of several splits; the most commonly used are EMNIST-Balanced and EMNIST-ByClass. The full dataset includes up to 814,255 $28\times 28$ grayscale images of 62 classes (digits 0–9, uppercase A–Z, and lowercase a–z). EMNIST introduces more structural diversity and class imbalance, especially in the cases of handwritten letters where individual writing styles cause considerable intra-class variation.
From a feature extraction standpoint, EMNIST provides a broad and challenging landscape. The similar morphology of certain characters (e.g., ‘O’ vs. ‘0’, ‘I’ vs. ‘l’) creates nontrivial inter-class similarity, making it a valuable testbed for evaluating fine-grained structural descriptors such as permutation entropy.
VI-D CIFAR-10 Dataset
CIFAR-10 is a benchmark dataset for general-purpose object classification in natural scenes [10]. It comprises 60,000 $32\times 32$ color images across 10 object categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. Unlike the previously discussed grayscale datasets, CIFAR-10 images are in full color and exhibit substantial variability in lighting, viewpoint, background, and object deformation.
The presence of color channels increases dimensionality and realism, but also introduces noise that can obscure structural patterns. While deep convolutional networks can achieve over 95% accuracy on CIFAR-10 [11], such performance is often driven by color and texture cues rather than purely structural features. For this reason, CIFAR-10 serves as a more difficult and less isolated setting for entropy-based feature extraction, requiring careful preprocessing or channel-wise analysis to focus on spatial complexity.
Comparative Analysis of the Dataset: The selected datasets—Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10—represent a diverse spectrum of visual complexity. Fashion-MNIST and KMNIST retain the simplicity of small grayscale images while introducing semantic and structural diversity. EMNIST increases the class space and handwriting variability, and CIFAR-10 brings in full-color imagery with real-world variability.
For our investigation of permutation entropy-based feature extraction, Fashion-MNIST, KMNIST, and EMNIST offer the most controlled yet challenging environments. Their grayscale format isolates structural and textural elements, making them ideal for analyzing spatial complexity. In contrast, CIFAR-10 introduces naturalistic noise and color-based variation, highlighting the limitations and potential adaptations needed when applying entropy-based techniques in more complex visual domains.
VII Image Classification Results and Discussion
VII-A Results on Fashion-MNIST Dataset
We apply the proposed feature extraction pipeline (Fig. 4) to the Fashion-MNIST dataset, that comprises $28\times 28$ grayscale images of fashion articles. Based on our method, each image is converted into a 780-dimensional feature vector, comprising entropy-based, correlation-based, shape-based and texture-based features (Fig. 5) as detailed below:
Row-wise PE: 28 rows → 28 features; Column-wise PE: 28 columns → 28 features; Row Correlation: 27 adjacent row pairs → 27 features; Column Correlation: 27 adjacent column pairs → 27 features; Diagonal PE: 21 diagonals ($k=-10$ to $10$) → 21 features; Anti-diagonal PE: 21 diagonals on horizontally flipped image → 21 features; Patch-wise PE: 169 overlapping patches of size $4\times 4$ with stride 2 → 169 features; HOG: 441 features using $4\times 4$ pixel cells and 9 orientation bins; LBP: 18-bin histogram with $P=16$, $R=2$ (uniform pattern) → 18 features.
TABLE II: Breakdown of Handcrafted Feature Types Used in Fashion-MNIST Dataset
Feature Type Number of Features
Row-wise PE 28
Column-wise PE 28
Diagonal PE 21
Anti-diagonal PE 21
Patch-wise PE 169
Row Correlation 27
Column Correlation 27
HOG 441
LBP 18
Total 780
Figure 5: Visualization of extracted features from a single Fashion-MNIST sample: (a) original 28×28 grayscale image; (b) patch-wise permutation-entropy (PE) heatmap computed over 4×4 windows with stride 2; (c) local binary pattern (LBP) texture map using P=16, R=2; (d) Histogram of Oriented Gradients (HOG) visualization with 4×4 pixels per cell; and (e) finally a normalized, concatenated feature vector used as input to the SVM classifier.
This 780-dimensional feature vector (Table II) is standardized and used as input to the classification models. Similar extraction procedures are applied to other datasets (KMNIST, EMNIST, CIFAR-10), with minor adjustments to patch count or image resolution as necessary. The resulting feature matrices had shapes $(60000,780)$ for training and $(10000,780)$ for testing. An SVM classifier with RBF kernel was tuned using 3-fold cross-validation over a grid of 12 $(C,\gamma)$ combinations, totaling 36 fits. The best hyperparameters obtained were $C=10$ and $\gamma=0.001$.The best cross-validation accuracy achieved during tuning was 90.44%, and the final test accuracy using the best SVM model was 91.15%. Table III presents a comparison of our results with previously reported SVM-based approaches on the Fashion-MNIST dataset that did not employ deep learning or handcrafted descriptors of comparable complexity.
TABLE III: Comparison with existing Classical ML-based methods on Fashion-MNIST (classical ML only)
Method Feature Type Accuracy
Logistic regression [13] Raw $28\times 28$ flattened 84.2%
Random forest [17] Raw pixels 87.3%
KNN Classifier[13] Raw pixels 85.4%
Linear SVM [13] Raw $28\times 28$ flattened 83.6%
RBF SVM [17] Raw $28\times 28$ flattened 89.7%
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 91.15%
The proposed method significantly improves over conventional SVM models trained on raw pixel intensities or texture-only features. By integrating entropy-based and correlation-based descriptors with traditional texture features like HOG and LBP, the model captures both statistical randomness and structural dependencies within the image. This comprehensive feature set enables better generalization and class separability, especially for visually similar classes in Fashion-MNIST. Compared to the best reported classical SVM accuracy of 89.7% [17], our method achieves a relative improvement of approximately 1.6% in absolute accuracy. This is a notable gain considering no deep learning models were used. Moreover, the method remains computationally feasible and inexpensive, with model training times ranging between 8–67 minutes depending on hyperparameters, which is acceptable for non-deep learning pipelines. This validates the effectiveness of the proposed handcrafted feature extraction method, and achieves superior performance compared to traditional SVM approaches reported in literature (Table III and Fig. 6).
VII-B Results on KMNIST Dataset
We apply the same feature extraction pipeline (Fig. 4) to the KMNIST dataset, which consists of grayscale images of Japanese Kuzushiji characters. Unlike Fashion-MNIST, KMNIST has more complex shapes and strokes, but our combination of permutation entropy (PE), correlation features, Histogram of Oriented Gradients (HOG), and Local Binary Patterns (LBP) successfully captured the intrinsic texture and structural information. The resulting feature matrices had shapes $(60000,780)$ for training and $(10000,780)$ for testing
We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=200$ and $\gamma=0.001$, yielding a cross-validation accuracy of 97.55%. On the held-out test set, the tuned RBF SVM achieved a test accuracy of 94.46%, outperforming previous classical machine learning SVM approaches on KMNIST that used raw pixels or simpler handcrafted features. Table IV summarizes the comparison.
The improvement can be attributed to the rich feature representation that combines multiple complementary descriptors. Permutation entropy captures the complexity and irregularity of pixel intensity patterns, correlation features encode spatial relationships, and HOG and LBP extract local shape and texture patterns.
TABLE IV: Comparison with existing Classical ML-based methods on KMNIST (classical ML only)
Method Feature Type Accuracy
4-Nearest Neighbour Baseline [20] Raw pixels 91.56%
PCA + 4-KNN [21] Raw pixels 93.98%
Tuned SVM [21] Raw pixels 92.82%
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 94.46%
Figure 6: Test accuracy comparison across benchmark datasets between previous classical machine learning methods and our approach (present study) that combines handcrafted feature extraction with SVM. Our method consistently outperforms traditional classical ML-based methods and feature descriptors, demonstrating its effectiveness and efficiency across Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10 dataset.
VII-C Results on EMNIST Dataset
We also tested our handcrafted feature extraction pipeline to the EMNIST Letters dataset, generating feature vectors of shape (124800, 780) for training and (20800, 780) for testing. Using a 3-fold cross-validation with grid search over a range of SVM hyperparameters, the best model was found with parameters $C=10$ and $\gamma=0.001$. The best cross-validation accuracy achieved was 92.83%, while the tuned SVM attained a test accuracy of 93.27%.
As shown in Table V, our method outperforms existing classical SVM pipelines applied to EMNIST datasets. The results demonstrate the effectiveness of combining multiple handcrafted features for the EMNIST handwritten character recognition task. The performance is consistent with or better than previously reported classical SVM-based methods on EMNIST, which typically utilize individual feature types such as raw pixels or HOG descriptors.
TABLE V: Comparison with existing Classical ML-based methods on EMNIST (classical ML only)
Method Feature Type Accuracy
Linear classifier [22] Raw pixels 55.78%
OPIUM[22] Raw pixels 85.15%
DWT-DCT + SVM [23] Extracted features 89.51%
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 93.27%
VII-D Results on CIFAR-10 Coloured Dataset
To apply our handcrafted feature pipeline to CIFAR-10, we extended our method to handle RGB color channels. Each image was split into its red, green, and blue channels. Features were extracted independently from each channel using the same feature extraction pipeline (Fig. 4). Each channel contributed a set of entropy and texture features, which were concatenated to form the final feature vector. In total, this resulted in a 1281-dimensional feature vector per image. We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=100$ and $\gamma=0.001$, yielding a cross-validation accuracy of 61.51%. On the held-out test set, the tuned SVM achieved an accuracy of 64.81%.
TABLE VI: Comparison with existing Classical ML-based methods on CIFAR-10 (classical ML only)
Method Feature Type Accuracy
SVM + PCA [24] Top 100 PCA components 40.2%
Linear SVM [25] 32 x 32 raw pixels 42.3%
SVM + Logistic Regression [26] SVM on raw pixels 57.8%
HOG + PCA + RBF SVM [27] Grayscaling and HOG 62.0%
Ours (PE + Corr + HOG + LBP) 1281 handcrafted RGB features 64.81%
As shown in Table VI, our method outperforms existing classical SVM pipelines applied to CIFAR-10. By combining entropy-based complexity measures (including patch-based permutation entropy), correlation structure, edge and texture descriptors (HOG and LBP), we capture richer information than standard approaches that rely solely on edge or intensity features.
Coates et al. [28] extracted a 4 000-dimensional feature vector by learning single-layer convolutional filters via K-means and simple pooling, achieving 79.6% accuracy on CIFAR-10. However, their approach incurs a significant offline training cost and yields features that are difficult to interpret. In contrast, our permutation-entropy + HOG + LBP pipeline produces a compact, lightweight feature descriptor that can be computed in a fully feed-forward manner without any filter learning, and whose individual components (entropy, gradient histograms, binary patterns) offer clear, intuitive interpretations. This interpretability and efficiency comes with a modest trade-off in accuracy, making our method particularly well-suited to applications with limited compute or a need for transparent feature representations.
VIII Conclusion
To conclude, we presented a unified handcrafted feature extraction pipeline that integrates diverse descriptors, including temporal-complexity measures such as permutation entropy, spatial correlations, edge- and texture-based features (e.g., HOG, LBP), to construct a rich and interpretable feature representation. By applying this approach to a suite of benchmark datasets—Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10—we demonstrated that shallow classifiers, particularly SVM, can achieve competitive performance without relying on deep learning architectures.
Our results show that combining complementary handcrafted features significantly outperforms traditional single-feature methods in the classical machine learning regime (Fig. 6). Notably, we achieve up to 91.15% test accuracy on Fashion-MNIST and 64.81% on CIFAR-10 using only interpretable, resource-efficient techniques. While still behind state-of-the-art deep neural networks, our method narrows the gap and offers a viable alternative for constrained environments where explainability, simplicity, or limited computational resources are priorities. These findings reinforce the relevance of classical approaches when enhanced with thoughtful feature engineering, and suggest that hybrid or complementary strategies may continue to play a role alongside deep learning in future vision systems.
In the future we plan to use these features as input representations for designing next-generation neural networks. Instead of feeding raw pixels into a deep model, we could build architectures that learn from more meaningful, structured inputs—like the features we extract here. This could lead to models that are faster to train, more accurate with less data, and easier to interpret. Therefore, combining the strengths of handcrafted features with the flexibility of deep learning could pave the way for next-generation computer vision systems that are both efficient and explainable.